# GATSY: Graph Attention Network for Music Artist Similarity thanks:

### v2

@@ Abstract
The artist similarity quest has become a crucial subject in social and scientific contexts, driven by the desire to enhance music discovery according to user preferences. Modern research solutions facilitate music discovery according to user tastes. However, defining similarity among artists remains challenging due to its inherently subjective nature, which can impact recommendation accuracy. This paper introduces GATSY, a novel recommendation system built upon graph attention networks and driven by a clusterized embedding of artists. The proposed framework leverages the graph topology of the input data to achieve outstanding performance results without relying heavily on hand-crafted features. This flexibility allows us the inclusion of fictitious artists within a music dataset, facilitating connections between previously unlinked artists and enabling diverse recommendations from various and heterogeneous sources. Experimental results prove the effectiveness of the proposed method with respect to state-of-the-art solutions while maintaining flexibility. The code to reproduce these experiments is available at https://anonymous.4open.science/r/GATSY-Music_Artist_Similarity-4807/README.md
@@ I INTRODUCTION
The artist similarity task involves finding a match between artists based on their shared aspects. However, there is no unambiguous way to define how two or more artists can be similar and the main characteristics that make them similar [1]. Similarity can be defined according to several points of view, such as cultural perspectives (e.g., provenance, historical context) or aspects more commonly related to the content of songs (e.g., common musical progressions, song keys). The ground truth for artist similarity does not rely solely on these individual features but rather on a combination of them. Such an approach is often implemented by extracting and elaborating music tracks. However, the resulting features cannot represent a subjective opinion concerning an artist categorization. Furthermore, digitalizing user tastes is always a critical task that inevitably affects a music recommendation system. In recent years, the research community has developed various artist similarity methods using Graph Neural Networks (GNNs) [2], owing to their natural ability to extract patterns from graph-structured data. Modeling artist relationships through graphs allows not only for the derivation of similarity scores but also for addressing challenges associated with long-tail artist recommendations [3, 4, 5]. However, generating the necessary artist data to train a GNN often requires information that is not readily available and may involve significant effort to collect, as highlighted in [6]. Furthermore, beyond the methodologies designed for data extraction, there are limited open-source datasets available for research purposes. In this work, we build on the idea of using GNNs for artist similarity and introduce a recommendation system that addresses these challenges. Our approach defines similarity via content-based and cultural-based aspects [6]. Experimental results show comparable accuracy to existing models while significantly reducing the number of parameters and offering a more efficient solution. The main contributions of this paper are as follows:
  • We introduce GATSY, a novel graph attention network-based model for artist similarity.
@@ II Related Works
Artist Similarity. Several methods recently attempted to address the artist similarity problem despite the difficulty in expressing an objective meaning of this similarity [1]. Natural language processing (NLP) methodologies found a role of relevance in this scenario by embedding a vast amount of text data describing artists and their songs. For instance, in [7], text features were used to define a knowledge base of artists. Also, tasks like the Named Entity Recognition (NER) [8] were applied to Wikipedia sources to obtain vector representations [9]. Besides relying on text attributes, artists have been also represented through their content-based features as in [10, 11]. In particular, in [10] it is also shown how heterogeneous features can outperform models with features-specific approaches.
@@
Graph Neural Networks. In the past decade, GNNs have contributed heavily to several areas of Deep Learning (DL), such as Recommendation Systems [12, 13, 14, 15], Natural Language Processing [16, 17, 18, 19, 20, 21], Computer Vision [22, 23, 24, 25], and Knowledge graphs [26, 27]. Generally, most of the GNNs follow the message-passing formalism [28], where nodes update iteratively their representations, after being trained on downstream tasks such as link prediction, and node or graph classification [2]. Due to the natural representation of music artists as a network, GNNs have found a place also in artist similarity applications.
@@
Artist Similarity with Graph Neural Networks. [6] is the seminal work that merges together artist similarity and GNNs. Here, is employed a graph of artists, where the connections are specified by user tastes, and each artist has an individual representation that is derived from track-level statistics about the loudness, dynamics and spectral shape of the signal. In this work is shown that through an unsupervised training strategy, the artist relationships and the individual features help effectively relate to non-previously connected artists. In an extension of this work, namely [3], have been proved that their approach allows also for long-tail artist recommendations, where some nodes present few connections w.r.t. those associated with popular artists. However, their approach heavily relies on the knowledge of input features for the artists, which are not always known in advance. Moreover, do not describe the effects of using more expressive GNNs, limiting their analysis to GraphSAGE [29] as a GNN candidate. GNNs have been employed to solve long-tail artist recommendations also in [4]. More recently in [5], were used more complex graphs, specifically, a multimodal representation is encoded through a heterogenous network, with information about the Lyrics, Audio, Artist. This approach formalizes the artist similarity computation as a Link Prediction task, where we estimate the probability that an edge exists between two nodes that resemble two artists. In this work, we also address artist similarity with GNNs and follow an approach similar to [6], since we compare on the same experimental set-up. We show that choosing a more expressive and efficient (in terms of parameters) architecture, is a way to overcome the need for hand-crafted artist features.
@@ III-A Addressing Artist Similarity with Graph Neural Networks
In [6], has been already proposed GraphSAGE layers [29] as GNN(,)($\cdot$,$\cdot$)( ⋅ , ⋅ ). In this work, we investigate other GNNs [30, 31], but we focus our experiments mainly on Graph Attention Network (GAT) [32], which shows superior performance w.r.t. the other GNNs for the task of artist similarity. GAT leverages the multi-headed self-attention mechanism, which helps to measure how much a neighbor attends to another node. This is a huge advantage when dealing with the heterophily of the graph at hand.
@@ III-B Loss Functions
The loss function minimizes the distance between the anchor and its positive while maximizing the distance from its negative. In this scenario, we defined positive samples as nodes in the same neighborhood and negative ones otherwise. We used the same approach illustrated in [34] to compute the triplets based on distance-weighted sampling. During the training phase, we adopted a minibatch setting as described in [29], while the triplets were chosen within the respective minibatch. This contrastive optimization approach is then coupled with a cross-entropy loss $\mathcal{L}_{\text{CE}}(\cdot)$ in the case of label availabilities. The cost function, when including the supervised objective is formalized as:
@@ III-C Model Architecture
The proposed GATSY framework is completely suited to the problem of music artist similarity. We derived an architecture based on two main blocks. The first one consists of a set of fully connected (FC) layers followed by a second block of graph convolution (GC) operations. This guarantees a limited model complexity while avoiding any over-smoothing. In particular, we use a set of 3 FC layers in the first block and 2 GC layers in the second one. This configuration involves at each transformation $256$ features, and it projects the $m$ input features to a $m’$ ’ vector space (where $m$ and $m’$ ’ are $2613$ and $256$ respectively). We also use it as the backbone architecture for the supervised learning task, including the artist-genre classification where we added on top of GATSY a set of 2 FC layers to predict the music genre. We use the exponential linear unit (ELU) [35] after each layer except the output one.
@@ IV Experimental Setup
In this Section, we provide the information concerning our experimental set-up. Our objective was to compare with [6], but their code is proprietary, and at the time of the experiments, the dataset was not freely available. We retrieved most of the dataset and kept a similar evaluation setting. In fact, our evaluation dataset is Olga, here we describe the modifications it underwent, and also the characteristics that we introduced. Later in this Section, we report also the models we included in our benchmark to compare against GATSY, which is the same denoted in [6].
@@ IV-A Original Dataset
We perform the model evaluation over Olga [6], which is one of the most popular datasets for music similarity [6]. Although its size is limited, it is considered one of the largest available datasets for this task. In this work, most of the experiments are based on the available samples of Olga. In fact, it was not possible to retrieve the whole dataset, according to the instructions provided in [6]. For this reason, we used a subgraph, whose description can be found in Table I. Besides the artist relations $\mathcal{A}$, the Olga dataset shows low-level audio features or input features, denoted as $\mathcal{X}$, representing track-level statistics about loudness, dynamics, and spectral shape of the signal. Moreover, these features include more abstract descriptors of rhythm and tonal information, such as the bpm and the average pitch class profile. The dataset is in the form of an undirected graph and we split it into training/validation/test sets in the percentages of 80/10/10 respectively.
@@ IV-B New Augmented Version of the Dataset
Driven by this need, we structured a new version of Olga containing genre labels for each artist in the graph. In the reduced Olga dataset, we attempted to retrieve the music genres through the MusicBrainz [^1] https://musicbrainz.org/ API. However, we were not able to retrieve labels for all the artists, so we based this retrieval phase also using NLP strategies.
@@
In this new version of the dataset, we explicitly list all the distinct $2842$ genres we found for the artists. However, in order to carry out an admissible and interpretable analysis, we wanted to consider only the $25$ most relevant and frequent ones, assuming that the excluded labels are associated sub-genres. Unfortunately, it was not always possible to have a unique musical genre for each artist, so disambiguation strategies were implemented as described below.
@@
  • Artists without an associated tag were identified with the most common musical genre among its neighbors in the graph.
@@
The new version of the dataset allows to extend the set of experiments, including not only an unsupervised regime but also a supervised one. In other words, we lead two sets of experiments: one is unsupervised and exploits only the loss function in Equation (5), while the second set of experiments encompasses also a supervised objective according to Equation (6). In both settings (supervised and unsupervised), the accuracy results have been tested with 10 different random seeds and show similar performance. Our goal is to improve the quality of our recommendation system by leveraging the new characterization of the nodes in the graph.
@@ IV-D State-of-the-Art Models for Comparisons
Conversely from [6], we also included batch normalization layers [42] after each transformation, except for the final one. This choice has been favoured by the advantages that it entails, mostly in terms of accuracy. The details on the hyperparameters are available in our codebase.
@@ V-A Artist Similarity
With the recent release of the full version of Olga, as specified in Table I, we have extended our experiments to the complete dataset. These results are presented in the Full Dataset box. We conducted these experiments using four different GNN architectures—GCN, GIN, GraphSAGE, and GAT—implementing them in accordance with the algorithm described in [6]. We performed extensive hyperparameter tuning for all models and trained them until convergence.
@@ V-B Ablation Study
To understand the key components of our method that contribute most to its overall performance, we refer to the ablation study presented in Figure 2 (reduced dataset) and Figure 3 (full dataset). They both present, across all the baselines, how performance drops depending on the type of input features used—whether random or hand-crafted. In Figure 2, in the FC and GraphSAGE models, we observe a similar phenomenon as reported in [6], despite the changes in the graph structure (see Table I), indicating that GraphSAGE is highly dependent on input features. The performance of the FC model, on the other hand, reinforces the idea that incorporating graph structure would enhance model performance. In the case of GATSY, Fig. 2 shows that the loss in performance is relatively low compared to the baselines. Specifically, the performance loss is 81.8% for FC, 74.89% for GraphSAGE, and only 11.38% for GATSY. Once, we had access to the full Olga dataset, we also repeated this very same ablation study, considering also the other GNN models. The results are reported in Figure 3. Also within the full dataset, we can assess that GATSY is the solution most resilient to random features, and GraphSAGE behaves as the one struggling the most. Also GCN and GIN seem resilient to input features as well, but their performance are slightly lower than GATSY.
@@
Figure 4: Visualization of the low-level features after dimensionality reduction (LEFT), and of the GATSY embedding after training on labeled Olga and dimensionality reduction (RIGHT). Here are shown the embeddings labelled among only the top-10 genre for visualization purposes.
TABLE III: Results gained from the experiment with the reduced dataset. We report the mean value with the related standard deviation, and the f1-score for artist genre classification. Accuracy refers to nDCG.
@@ V-C Supervised Learning Experiments
In this Section, we describe the performance achieved in the supervised experiments. Here, we rely on the labelled dataset, where the number of nodes is further reduced (see Table I). We structure new splits and provide a new baseline for the unsupervised case (see GATSY in Table III). This way, we can show the performance variation of the model when the objective becomes the one in Equation (6). In Table III we show the performance of the genre-aware model. We confirm that the $\text{nDCG}_{200}$ is more expressive as an unsupervised metric. Indeed, the unsupervised approach is always outperforming the supervised experiments in terms of accuracy. Despite this, the supervised trained GATSY becomes capable of predicting the artist labels, which means that it is more genre-aware than the unsupervised trained GATSY. Furthermore, we expect that achieving a high f1-score is quite challenging as associating an artist with a unique genre is rather unfeasible. This is also the reason why the supervised loss struggles to help the unsupervised one. For instance, in the Olga graph, rarely the neighborhood is identified by a unique class, as corroborated by our discussion on the homophily level of Olga (see Section IV-C). In this sense, the supervised and unsupervised objectives may be colliding whenever two neighbors of different classes attract themselves to minimize $\mathcal{L}_{\text{tri}}$, and also repulsing to minimize $\mathcal{L}_{\text{CE}}$. These conclusions allow us to answer negatively to RQ4.
In Fig. 4, we show the low-level features associated with the music genre, and the clusters we obtain when GATSY is trained both on $\mathcal{L}_{\text{tri}}$ and $\mathcal{L}_{\text{CE}}$. Here the embeddings organize themselves in the space driven by the supervision signal associated with the music genre. The clusters are identified by the ground-truth values in $\mathcal{Y}$.

### v3

@@ Abstract
The artist similarity quest has become a crucial subject in social and scientific contexts, driven by the desire to enhance music discovery according to user preferences. Modern research solutions facilitate music discovery according to user tastes. However, defining similarity among artists remains challenging due to its inherently subjective nature, which can impact recommendation accuracy. This paper introduces GATSY, a novel recommendation system built upon graph attention networks and driven by a clusterized embedding of artists. The proposed framework leverages the graph topology of the input data to achieve outstanding performance results without relying heavily on hand-crafted features. This flexibility allows us to include fictitious artists within a music dataset, facilitating connections between previously unlinked artists and enabling diverse recommendations from various and heterogeneous sources. Experimental results prove the effectiveness of the proposed method with respect to state-of-the-art solutions while maintaining flexibility. The code to reproduce these experiments is available at https://github.com/difra100/GATSY-Music_Artist_Similarity.
@@
11footnotetext: Department of Computer Science, Control and Management Engineering, Sapienza University of Rome, Rome, Italy.22footnotetext: Institute of Information Science and Technologies ”Alessandro Faedo” - ISTI-CNR, Pisa, Italy.33footnotetext: Department of Computer Science, Sapienza University of Rome, Rome, Italy.55footnotetext: Department of Information Engineering, Electronics and Telecommunications, Sapienza University of Rome, Rome, Italy.§§footnotetext: Corresponding author, difrancesco@diag.uniroma1.it.
@@ I INTRODUCTION
The artist similarity task involves finding a match between artists based on their shared aspects. However, there is no unambiguous way to define how two or more artists can be similar and the main characteristics that make them similar [1]. Similarity can be defined according to several points of view, such as cultural perspectives (e.g., provenance, historical context) or aspects more commonly related to the content of songs (e.g., common musical progressions, song keys). The ground truth for artist similarity does not rely solely on these individual features but rather on a combination of them. Such an approach is often implemented by extracting and elaborating music tracks. However, the resulting features cannot represent a subjective opinion concerning an artist’s categorization. Furthermore, digitalizing user tastes is always a critical task that inevitably affects a music recommendation system. In recent years, the research community has developed various artist similarity methods using Graph Neural Networks (GNNs) [2], owing to their natural ability to extract patterns from graph-structured data. Modeling artist relationships through graphs allows not only for the derivation of similarity scores but also for addressing challenges associated with long-tail artist recommendations [3, 4, 5]. However, generating the necessary artist data to train a GNN often requires information that is not readily available and may involve significant effort to collect, as highlighted in [6]. Furthermore, beyond the methodologies designed for data extraction, there are limited open-source datasets available for research purposes. In this work, we build on the idea of using GNNs for artist similarity and introduce a recommendation system that addresses these challenges. Our approach defines similarity via content-based and cultural-based aspects [6]. Experimental results show comparable accuracy to existing models while significantly reducing the number of parameters and offering a more efficient solution. The main contributions of this paper are as follows:
  • We introduce GATSY (Graph ATtention network for music artist SimilaritY), a novel graph attention network-based model for artist similarity.
@@ II Related Works
Artist Similarity. Several methods recently attempted to address the artist similarity problem despite the difficulty in expressing an objective meaning of this similarity [1]. Natural language processing (NLP) methodologies found a role of relevance in this scenario by embedding a vast amount of text data describing artists and their songs. For instance, in [7], text features were used to define a knowledge base of artists. Also, tasks like the Named Entity Recognition (NER) [8] were applied to Wikipedia sources to obtain vector representations [9]. Besides relying on text attributes, artists have also been represented through their content-based features as in [10, 11]. In particular, in [10] it is also shown how heterogeneous features can outperform models with feature-specific approaches.
@@
Graph Neural Networks. In the past decade, GNNs have contributed heavily to several areas of Deep Learning (DL), such as Recommendation Systems [12, 13], Natural Language Processing [14, 15, 16, 17, 18, 19], Computer Vision [20, 21, 22, 23], and Knowledge graphs [24, 25]. Generally, most of the GNNs follow the message-passing formalism [26], where nodes update iteratively their representations, after being trained on downstream tasks such as link prediction, and node or graph classification [2]. Due to the natural representation of music artists as a network, GNNs have also found a place in artist similarity applications.
@@
Artist Similarity with Graph Neural Networks. [6] is the seminal work that merges together artist similarity and GNNs. Here, a graph of artists is employed, where user tastes specify the connections, and each artist has an individual representation that is derived from track-level statistics about the loudness, dynamics, and spectral shape of the signal. This work shows that through an unsupervised training strategy, the artist relationships and the individual features help relate effectively to non-previously connected artists. In an extension of this work, namely [3], it has been proved that their approach also allows for long-tail artist recommendations, where some nodes present few connections w.r.t. those associated with popular artists. However, their approach heavily relies on the knowledge of input features for the artists, which are not always known in advance. Moreover, do not describe the effects of using more expressive GNNs, limiting their analysis to GraphSAGE [27] as a GNN candidate. GNNs have been employed to solve long-tail artist recommendations also in [4]. More recently, in [5], multimodal representations and heterogenous networks are used to capture more intricate artist relationships, considering explicitly also Lyrics, and Audio. This approach formalizes the artist similarity computation as a Link Prediction task, where we estimate the probability that an edge exists between two nodes that resemble two artists. In this work, we also address artist similarity with GNNs and follow an approach similar to [6], since we compare on the same experimental setup. We show that choosing a more expressive and efficient (in terms of parameters) architecture is a way to overcome the need for hand-crafted artist features.
@@ III-A Addressing Artist Similarity with Graph Neural Networks
In [6], GraphSAGE layers [27] have already been proposed as GNN(,)($\cdot$,$\cdot$)( ⋅ , ⋅ ). In this work, we investigate other GNNs [28, 29], but we focus our experiments mainly on Graph Attention Network (GAT) [30], which shows superior performance w.r.t. the other GNNs for the task of artist similarity. GAT leverages the multi-headed self-attention mechanism, which helps to measure how much a neighbor attends to another node. This is a huge advantage when dealing with the heterophily of the graph at hand.
@@ III-B Loss Functions
The loss function minimizes the distance between the anchor and its positive while maximizing the distance from its negative. In this scenario, we defined positive samples as nodes in the same neighborhood and negative ones otherwise. We used the same approach illustrated in [32] to compute the triplets based on distance-weighted sampling. During the training phase, we adopted a minibatch setting as described in [27], while the triplets were chosen within the respective minibatch. This contrastive optimization approach is then coupled with a cross-entropy loss $\mathcal{L}_{\text{CE}}(\cdot)$ in the case of label availabilities. The cost function, when including the supervised objective, is formalized as:
@@ III-C Model Architecture
The proposed GATSY framework is completely suited to the problem of music artist similarity. We derived an architecture based on two main blocks. The first one consists of a set of fully connected (FC) layers followed by a second block of graph convolution (GC) operations. This guarantees a limited model complexity while avoiding any over-smoothing. In particular, we use a set of 3 FC layers in the first block and 2 GC layers in the second one. This configuration involves, at each transformation $256$ features, and it projects the $m$ input features to a $m’$ ’ vector space (where $m$ and $m’$ ’ are $2613$ and $256$ respectively). We also use it as the backbone architecture for the supervised learning task, including the artist-genre classification, where we added on top of GATSY a set of 2 FC layers to predict the music genre. We use the exponential linear unit (ELU) [33] after each layer except the output one, similarly to [6].
@@ IV Experimental Setup
In this Section, we provide the information concerning our experimental setup. Our objective was to compare with [6], but their code is proprietary, and at the time of the experiments, the dataset was not freely available. We retrieved most of the dataset and kept a similar evaluation setting. In fact, our evaluation dataset is Olga, which is one of the most popular datasets for music similarity [6].,here we describe the modifications it underwent, and also the characteristics that we introduced. Later in this Section, we also report the models we included in our benchmark to compare against GATSY, which is the same as denoted in [6].
@@ IV-A Original Dataset
We perform the model evaluation over Olga [6], which is one of the most popular datasets for music similarity [6]. Although its size is limited, it is considered one of the largest available datasets for this task. In this work, most of the experiments are based on the available samples of Olga. In fact, it was not possible to retrieve the whole dataset, according to the instructions provided in [6]. For this reason, we used a subgraph, whose description can be found in Table I. Besides the artist relations $\mathcal{A}$, the Olga dataset shows low-level audio features or input features, denoted as $\mathcal{X}$, representing track-level statistics about loudness, dynamics, and spectral shape of the signal. Moreover, these features include more abstract descriptors of rhythm and tonal information, such as the bpm and the average pitch class profile. The dataset is in the form of an undirected graph and we split it into training/validation/test sets in the percentages of 80/10/10, respectively.
@@ IV-B New Augmented Version of the Dataset
Driven by this need, we structured a new version of Olga containing genre labels for each artist in the graph. In the reduced Olga dataset, we attempted to retrieve the music genres through the MusicBrainz [^1] https://musicbrainz.org/ API. However, we were not able to retrieve labels for all the artists, so we also based this retrieval phase using NLP strategies.
@@
In this new version of the dataset, we explicitly list all the distinct $2842$ genres we found for the artists. However, to carry out an admissible and interpretable analysis, we wanted to consider only the $25$ most relevant and frequent ones, assuming that the excluded labels are related sub-genres. Unfortunately, it was not always possible to have a unique musical genre for each artist, so disambiguation strategies were implemented as described below.
@@
  • Artists without an associated tag were identified with the most common musical genre among their neighbors in the graph.
@@
The new version of the dataset allows for extending the set of experiments, including not only an unsupervised regime but also a supervised one. In other words, we lead two sets of experiments: one is unsupervised and exploits only the loss function in Equation (5), while the second set of experiments encompasses also a supervised objective according to Equation (6). In both settings (supervised and unsupervised), the accuracy results have been tested with 10 different random seeds and show similar performance. Our goal is to improve the quality of our recommendation system by leveraging the new characterization of the nodes in the graph.
@@ IV-D State-of-the-Art Models for Comparisons
Conversely, from [6], we also included batch normalization layers [40] after each transformation, except for the final one. This choice has been favoured by the advantages that it entails, mostly in terms of accuracy. The details on the hyperparameters are available in our codebase.
@@ V-A Artist Similarity
With the recent release of the full version of Olga, as specified in Table I, we have extended our experiments to the complete dataset. These results are presented in the Full Dataset box. We conducted these experiments using four different GNN architectures—GCN, GIN, GraphSAGE, and GAT—implementing them following the algorithm described in [6]. We performed extensive hyperparameter tuning for all models and trained them until convergence.
@@ V-B Ablation Study
To understand the key components of our method that contribute most to its overall performance, we refer to the ablation study presented in Figure 2 (reduced dataset) and Figure 3 (full dataset). They both present, across all the baselines, how performance drops depending on the type of input features used—whether random or hand-crafted. In Figure 2, in the FC and GraphSAGE models, we observe a similar phenomenon as reported in [6], despite the changes in the graph structure (see Table I), indicating that GraphSAGE is highly dependent on input features. The performance of the FC model, on the other hand, reinforces the idea that incorporating graph structure would enhance model performance. In the case of GATSY, Fig. 2 shows that the loss in performance is relatively low compared to the baselines. Specifically, the performance loss is 81.8% for FC, 74.89% for GraphSAGE, and only 11.38% for GATSY. Once, we had access to the full Olga dataset, we also repeated this very same ablation study, considering also the other GNN models. The results are reported in Figure 3. Also within the full dataset, we can assess that GATSY is the solution most resilient to random features, and GraphSAGE behaves as the one struggling the most. Also GCN and GIN seem resilient to input features as well, but their performance is slightly lower than GATSY.
@@
Figure 4: Visualization of the low-level features after dimensionality reduction (LEFT), and of the GATSY embedding after training on labeled Olga and dimensionality reduction (RIGHT). Here are shown the embeddings labelled among only the top 10 genres for visualization purposes.
TABLE III: Results gained from the experiment with the reduced dataset. We report the mean value with the related standard deviation and the F1-score for artist genre classification. Accuracy refers to nDCG.
@@ V-C Supervised Learning Experiments
In this Section, we describe the performance achieved in the supervised experiments. Here, we rely on the labelled dataset, where the number of nodes is further reduced (see Table I). We structure new splits and provide a new baseline for the unsupervised case (see GATSY in Table III). This way, we can show the performance variation of the model when the objective becomes the one in Equation (6). In Table III we show the performance of the genre-aware model. We confirm that the $\text{nDCG}_{200}$ is more expressive as an unsupervised metric. Indeed, the unsupervised approach is always outperforming the supervised experiments in terms of accuracy. Despite this, the supervised trained GATSY becomes capable of predicting the artist labels, which means that it is more genre-aware than the unsupervised trained GATSY. Furthermore, we expect that achieving a high F1-score is quite challenging, as associating an artist with a unique genre is rather unfeasible. This is also the reason why the supervised loss struggles to help the unsupervised one. For instance, in the Olga graph, the neighborhood is identified by a unique class, as corroborated by our discussion on the homophily level of Olga (see Section IV-C). In this sense, the supervised and unsupervised objectives may be colliding whenever two neighbors of different classes attract themselves to minimize $\mathcal{L}_{\text{tri}}$, and also repulse to minimize $\mathcal{L}_{\text{CE}}$. These conclusions allow us to answer negatively to RQ4.
In Fig. 4, we show the low-level features associated with the music genre, and the clusters we obtain when GATSY is trained both on $\mathcal{L}_{\text{tri}}$ and $\mathcal{L}_{\text{CE}}$. Here, the embeddings organize themselves in the space driven by the supervision signal associated with the music genre. The clusters are identified by the ground-truth values in $\mathcal{Y}$.